# Feature-Enhanced Neural Collaborative Reasoning with Dense Representation for Recommendation

### v1

@@
Abstract
In recent years, various studies have focused on designing logical reasoning enhanced recommendation methods. However, previous studies either combine dense vector representation with relatively simple logical operators, or learn logical rule sets with one-hot encoding features. In this paper, we aim to simultaneously learn complex logical rules and utilize dense feature representations, improving model performance in both reasoning and explaining. We propose a novel model named Feature-Enhanced Neural Collaborative Reasoning (FENCR) in an end-to-end manner. Firstly, we construct dense feature embeddings with a selection module to automatically generate distinct user and item feature groups. Then, neural logical modules are designed to conduct logical operations (i.e., conjunction and disjunction) over the selected feature groups, by which logical rules are generated. Self-supervised logical axiom regularizers are accordingly introduced in the loss function to augment the logical properties of the modules. Finally, recommendation results are derived by personalized aggregation of multiple generated logical rules. Comparative experiments on two large public datasets show our model outperforms state-of-the-art logic-enhanced recommendation models. Further analyses also showcase the explainability of the model facilitated by derived rules.
Index Terms:
collaborative reasoning, interpretable modeling, explainable recommendation, rule learning
I Introduction
@@
Explainability has always been a topic of great concern within the field of recommendation [ExplainSurveyYongfengZhang, chen2022measuring, sun2020dual, sun2021unsupervised]. Researchers explore various methods to help users understand why recommendation systems give certain results. Among these methods, the text-based explanation emerges as a kind of prominent and widely-used approach [DBLP:GenerateLanguage, li2021personalizedTransformer, li2022personalized]. Through explanation text, systems can present the reason for recommendations to users in natural language, thereby increasing user trust and comprehensibility of the given results.
@@
An effective evaluation should ensure the explanations truly resonate with users and meet their expectations. However, with advanced approaches developed for generating explanatory text, assessing their quality remains an issue that has yet to be adequately resolved. Existing methods for evaluation can be categorized into three main types: self-report, third-party annotations, and reference-based metrics. Conducting user studies to obtain self-reported feedback can most accurately reflect user experience. This approach requires evaluations to be recorded alongside the recommendations and is, therefore, difficult to obtain and use in public datasets. Third-party annotations can reflect human feedback and are relatively accessible. Still, manual labeling is expensive, time-consuming, and lack of scalability. Reference-based metrics assess the quality of the target text based on the reference text and offer a standard assessment that is relatively easy to acquire. Common metrics such as BLEU [DBLP:bleu2022ACL], and ROUGE [Rouge2004Lin] calculate the similarity between the generated text and the reference text. However, there may not exist reference texts for the scenario (many use reviews as a substitute). In addition, these textual similarities metrics may not ideally reflect user perception of recommendation explanation [ExpScore2022WWW, stopbleu]. These limitations highlight the need to develop evaluation methodologies that are in-line with human experience, easy to acquire, and reproducible.
@@
Recently, the development of large language models (LLM) sheds new light on the evaluation of various neural language generation (NLG) tasks [kocmi2023LLMEvaluateTransaltion, wang2023automatedpersonaliztion, wang2023chatgpt, fu2023gptscore]. Considering that LLMs can follow human instructions and their language modeling ability, LLMs can make adequate evaluations under reference-free settings with appropriate prompt [GEval2023liu, he2024annollm]. LLM offers an appealing solution for evaluating the quality of recommendation explanations since it is efficient (lower cost than manual labeling) and widely applicable (almost no dataset limitation). In a study on the design of explainable recommendation method [lei2023recexplainer], researchers introduce LLMs to evaluate the quality of explanations generated. However, we state that successes in the evaluation of general NLG tasks can not be migrated to the evaluation of explainable recommendations without verification. Unlike those NLG tasks that are already been examined, evaluating recommendation explanation text is more sophisticated. The reason is that measuring the quality of recommendations involves a group of diverse goals, e.g. persuasiveness, transparency [ConflictGoalsOfExplanations], etc. Additionally, a large portion of these goals are related to the subjective perception of users. All these factors add to the difficulty of assessing the quality of recommendation explanation texts. These considerations underscore the need to further explore the feasibility of leveraging LLM as a potential solution for text-based explainable recommendation evaluation.
@@
To answer these research questions, we use data from a user study in previous work [UserPerceptionTois], including the users’ ratings on 4 aspects of the provided explanatory texts, as well as the self-explanation text of the users for the recommended movies. Our study is based on the premise that the user receiving the recommendation is the ground-truth evaluator for the quality of the explanation. To study the reliability of LLM annotations, we additionally collect third-party annotations and LLM annotations under human instructions. To comprehensively compare between evaluators, we design and apply a 3-level meta-evaluation strategy to measure the correlation between evaluator annotations and user labels on different aspects. We compare the evaluation accuracy of LLM evaluators with third-party annotations and commonly used reference-based metrics, i.e., BLEU [DBLP:bleu2022ACL] and ROUGE [Rouge2004Lin].
@@
II Experimental Setup
Figure 1: Traditional evaluation approaches vs. utilizing LLMs for evaluations.
II-A Problem Formulation
@@
II-B Data Construction
II-B1 Data overview
[…]
UserPerceptionTois create a movie recommendation platform. It first captures user preferences, and then gives personalized recommendations along with text-formed explanations generated by a series of systems. We utilize two parts of the data: 1) self-explanations written by users, 2) 1-5 ratings of users to explanations generated by different methods in terms of 4 aspects, which are persuasiveness, transparency, accuracy, and satisfaction.
Formally, 39 participants are included in $\mathcal{U}$ and $\mathcal{I}$ are movies from Movielens Latest dataset [^1] https://grouplens.org/datasets/movielens/latest/. $\mathcal{I}_{u}$ denotes the itemset recommended to the user $u$, which includes the top-8 movies calculated by BiasedMF [RendleMF2012]. $\mathcal{G}$ includes a series of systems used to generate the explanatory texts. Most of these methods are template-based methods, e.g. user-based method, which generates explanations in the form of: “[N%] of users who share similar watching tastes with you like [MOVIE TITLE] after watching it.” $\mathcal{G}$ also includes the system that directly generates the complete natural language sentence, i.e., peer-explanation written by others. The data includes the self-explanation for each user-item pair. Since they are written by users themselves, there is no evaluation rating attached.
@@ II-B1 Data overview
II-B2 Evaluations from users
@@
II-B3 Evaluation from third-party annotators
@@
II-B4 Evaluation from Quantitative metrics
We utilize BLEU [DBLP:bleu2022ACL] and ROUGE [Rouge2004Lin] to evaluate the explanatory text, taking the self-explanations written by users as reference texts. The evaluation score $\mathbf{\hat{s}}_{r}$ is a float that measures text similarity.
II-C LLM as Evaluator
II-C1 LLM Evaluator Construction
We utilize pre-trained LLMs to provide annotations for each explanation text $E$. The large language model receives the movie name and the corresponding explanation text, accompanied by a prompt $\mathcal{P}$ to describe the context. Although $E$ is in Chinese as introduced in Section II-B1, we apply $\mathcal{P}$ in English to ensure a standardized prompt design. The LLM evaluation for each explanatory text $\mathbf{\hat{s}}_{LLM(u,i,E)}$ is a one-dimensional vector of length 4, with elements being integers between 1 to 5. Our experiments are based on item-wise evaluation, where only one text entry is evaluated at a time. This approach aligns with the setup in the user study.
@@ II-C1 LLM Evaluator Construction
Figure 2: The outline of evaluation prompt templates applied in our study.
II-C2 Prompt Construction
@@
Single-Aspect vs. Multiple-Aspect. The difference between single and multiple-aspect prompts lies in the Aspect part. A multiple-aspect prompt instructs the model to evaluate all four aspects concurrently, while a single-aspect prompt focuses on one aspect at a time. Utilizing single-aspect prompts necessitates more times of interactions with LLM compared to multiple-aspect prompts. Experimental results under the two settings are presented and compared in RQ1, Section II-D.
{algorithm}
Comparison of personalized and non-personalized one-shot learning process
PersonalizedOneShot each user $u$in $\mathcal{U}$Randomly choose item $i_{0}$from $\mathcal{I}_{u}$ollect user evaluations on explanatory text $E=g(u,i_{0})$as $\mathbf{s_{0}}_{g}$for each $g$in $\mathcal{G}$each movie $i$in $\mathcal{I}_{u}$, each $g$in $\mathcal{G}$$Prompt\leftarrow$ConstructPrompt($i$, $g(u,i)$, $\mathbf{s_{0}}_{g}$) $Result\leftarrow$LLM($Prompt$) Parse $Result$and process the answer Non-PersonalizedOneShot Randomly choose user-item pair $(u_{0},i_{0})$Collect user evaluations on explanatory text $E=g(u_{0},i_{0})$as $\mathbf{s_{0}}_{g}$for each $g$in $\mathcal{G}$each user $u$in $\mathcal{U}$,each movie $i$in $\mathcal{I}_{u}$, each $g$in $\mathcal{G}$$Prompt\leftarrow$ConstructPrompt($i$, $g(u,i)$, $\mathbf{s_{0}}_{g}$) $Result\leftarrow$LLM($Prompt$) Parse $Result$and process the answer
[…]
@@
To explore whether humans and LLM can collaborate on evaluation, we investigate the impact of utilizing human labels as contextual information in LLM evaluation. Our primary focus is on one-shot learning and includes one human label in the prompt. To enhance the LLM’s ability to learn personalized preferences, we employ personalized one-shot learning. This approach involves providing the scoring example from the same user for the target data. Formally, for the user $u$, item $i$ and explanation $E=g(u,i)$, the personalized example is the scoring given to $E^{\prime}=g(u,i_{0})$, which is the explanation generated by the same system $g$ and the same user $u$ on a randomly chosen movie item $i_{0}$. The methods help the prompt $\mathcal{P}$ incorporate personal information. However, a limitation is that it still requires collecting real user feedback, which is expansive and sometimes impractical. Hence, we also investigate non-personalized one-shot prompts. In this setting, all examples are from randomly selected $(u_{0},i_{0})$ pair. That is, the example can come from another user and are easier to collect in practice. Formally, for the user $u$, item $i$ and explanation $E=g(u,i)$, the non-personalized example is the scoring given to $E^{\prime}=g(u_{0},i_{0})$. We summarize the personalized and non-personalized one-shot learning procedures in Algorithm 1 to better illustrate the example selection process and their differences. Experimental results under different example settings are presented and compared in RQ2, Section LABEL:sec:RQ2.
II-D Three-Level Meta Evaluation
Good evaluation requires that evaluation efforts themselves be evaluated [stufflebeam1974metaevaluation]. Employing the suitable meta-evaluation method $h(\cdot)$ is crucial for thoroughly examining the evaluation procedure. Correlation coefficients, such as Pearson($r$), Spearman($\rho$), and Kendall($\tau$) coefficients, have served as a widely used metric for gauging the similarity in trends between two arrays of ratings or scores [liu2023calibrating]. In NLG, the strategy of meta-evaluation can be divided into two levels: Dataset-Level and Sample-Level [liu2023calibrating]. Formally, given a dataset $\mathcal{D}$ consisting of a set of source data $\mathcal{X}=(x_{1},x_{2},....,x_{|\mathcal{X}|})$ and generation methods $\mathcal{G}=(g_{1}(\cdot),g_{2}(\cdot),...,g_{|\mathcal{G}|}(\cdot))$ and correlation metric $r(\cdot)$ dataset-level and sample-level meta evaluation are expressed as:
@@ II-D Three-Level Meta Evaluation
We use a concrete example to illustrate the possible variation of the same evaluation metric at different levels. For instance, our experimental results in Section II-D found that BLEU-4 scoring results correlate poorly or even negatively with ground-truth labels at the Dataset-Level, whereas correlations improve at User-Level and Pair-Level. This discrepancy arises because BLEU-4 computes the token similarity between the target text and the reference text. Some reference texts may contain more commonly used token combinations than others, potentially inflating the BLEU-4 score. This introduces a bias into the evaluation process that is unrelated to user perceptions, thereby impacting the Dataset-Level correlation. Full experimental results and analyses can be found in Section II-D.
TABLE I: The 3-level Pearson correlation between the results given by the evaluator and the ground-truth label given by users. Bold fonts denote the best results among all tested evaluators and the underlines show the second-best results.

### v2

# Large Language Models as Evaluators for Recommendation Explanations

@@
Abstract.
The explainability of recommender systems has attracted significant attention in academia and industry. Many efforts have been made for explainable recommendations, yet evaluating the quality of the explanations remains a challenging and unresolved issue. In recent years, leveraging LLMs as evaluators presents a promising avenue in Natural Language Processing tasks (e.g., sentiment classification, information extraction), as they perform strong capabilities in instruction following and common-sense reasoning. However, evaluating recommendation explanatory texts is different from these NLG tasks, as its criteria are related to human perceptions and are usually subjective.
In this paper, we investigate whether LLMs can serve as evaluators of recommendation explanations. To answer the question, we utilize real user feedback on explanations given from previous work and additionally collect third-party annotations and LLM evaluations. We design and apply a 3-level meta-evaluation strategy to measure the correlation between evaluator labels and the ground truth provided by users. Our experiments reveal that LLMs, such as GPT4, can provide comparable evaluations with appropriate prompts and settings. We also provide further insights into combining human labels with the LLM evaluation process and utilizing ensembles of multiple heterogeneous LLM evaluators to enhance the accuracy and stability of evaluations. Our study verifies that utilizing LLMs as evaluators can be an accurate, reproducible and cost-effective solution for evaluating recommendation explanation texts. Our code is available here [^1] https://github.com/Xiaoyu-SZ/LLMasEvaluator.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
Explainability has always been a topic of great concern within the field of recommendation (Zhang and Chen, 2020; Chen et al., 2022; Sun et al., 2020, 2021). Researchers explore various methods to help users understand why recommendation systems give certain results. Among these methods, the text-based explanation emerges as a kind of prominent and widely-used approach (Chen et al., 2021a; Li et al., 2021, 2022). Through explanation text, systems can present the reason for recommendations to users in natural language, thereby increasing user trust and comprehensibility of the given results.
@@
An effective evaluation should ensure the explanations truly resonate with users and meet their expectations. However, with advanced approaches developed for generating explanatory text, assessing their quality remains an issue that has yet to be adequately resolved. Existing methods for evaluation can be categorized into three main types: self-report, third-party annotations, and reference-based metrics. Conducting user studies to obtain self-reported feedback can most accurately reflect user experience. This approach requires evaluations to be recorded alongside the recommendations and is, therefore, difficult to obtain and use in public datasets. Third-party annotations can reflect human feedback and are relatively accessible. Still, manual labeling is expensive, time-consuming, and lack of scalability. Reference-based metrics assess the quality of the target text based on the reference text and offer a standard assessment that is relatively easy to acquire. Common metrics such as BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004) calculate the similarity between the generated text and the reference text. However, there may not exist reference texts for the scenario (many use reviews as a substitute). In addition, these textual similarities metrics may not ideally reflect user perception of recommendation explanation (Wen et al., 2022; Freitag et al., 2022). These limitations highlight the need to develop evaluation methodologies that are in-line with human experience, easy to acquire, and reproducible.
@@
Recently, the development of large language models (LLM) sheds new light on the evaluation of various neural language generation (NLG) tasks (Kocmi and Federmann, 2023; Wang et al., 2023a, b; Fu et al., 2023). Considering that LLMs can follow human instructions and their language modeling ability, LLMs can make adequate evaluations under reference-free settings with appropriate prompt (Liu et al., 2023a; He et al., 2024). LLM offers an appealing solution for evaluating the quality of recommendation explanations since it is efficient (lower cost than manual labeling) and widely applicable (almost no dataset limitation). In a study on the design of explainable recommendation method (Lei et al., 2023), researchers introduce LLMs to evaluate the quality of explanations generated. However, we state that successes in the evaluation of general NLG tasks can not be migrated to the evaluation of explainable recommendations without verification. Unlike those NLG tasks that are already been examined, evaluating recommendation explanation text is more sophisticated. The reason is that measuring the quality of recommendations involves a group of diverse goals, e.g. persuasiveness, transparency (Balog and Radlinski, 2020), etc. Additionally, a large portion of these goals are related to the subjective perception of users. All these factors add to the difficulty of assessing the quality of recommendation explanation texts. These considerations underscore the need to further explore the feasibility of leveraging LLM as a potential solution for text-based explainable recommendation evaluation.
@@
To answer these research questions, we use data from a user study in previous work (Lu et al., 2023), including the users’ ratings on 4 aspects of the provided explanatory texts, as well as the self-explanation text of the users for the recommended movies. Our study is based on the premise that the user receiving the recommendation is the ground-truth evaluator for the quality of the explanation. To study the reliability of LLM annotations, we additionally collect third-party annotations and LLM annotations under human instructions. To comprehensively compare between evaluators, we design and apply a 3-level meta-evaluation strategy to measure the correlation between evaluator annotations and user labels on different aspects. We compare the evaluation accuracy of LLM evaluators with third-party annotations and commonly used reference-based metrics, i.e., BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004).
@@
2. Experimental Setup
Figure 1. Traditional evaluation approaches vs. utilizing LLMs for evaluations.
2.1. Problem Formulation
@@
2.2. Data Construction
2.2.1. Data overview
Lu et al. (2023) create a movie recommendation platform. It first captures user preferences, and then gives personalized recommendations along with text-formed explanations generated by a series of systems. We utilize two parts of the data: 1) self-explanations written by users, 2) 1-5 ratings of users to explanations generated by different methods in terms of 4 aspects, which are persuasiveness, transparency, accuracy, and satisfaction.
Formally, 39 participants are included in $\mathcal{U}$ and $\mathcal{I}$ are movies from Movielens Latest dataset [^2] https://grouplens.org/datasets/movielens/latest/. $\mathcal{I}_{u}$ denotes the itemset recommended to the user $u$, which includes the top-8 movies calculated by BiasedMF (Rendle et al., 2012). $\mathcal{G}$ includes a series of systems used to generate the explanatory texts. Most of these methods are template-based methods, e.g. user-based method, which generates explanations in the form of: “[N%] of users who share similar watching tastes with you like [MOVIE TITLE] after watching it.” $\mathcal{G}$ also includes the system that directly generates the complete natural language sentence, i.e., peer-explanation written by others. The data includes the self-explanation for each user-item pair. Since they are written by users themselves, there is no evaluation rating attached.
@@ 2.2.1. Data overview
2.2.2. Evaluations from users
@@
2.2.3. Evaluation from third-party annotators
@@
2.2.4. Evaluation from Quantitative metrics
We utilize BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004) to evaluate the explanatory text, taking the self-explanations written by users as reference texts. The evaluation score $\mathbf{\hat{s}}_{r}$ is a float that measures text similarity.
2.3. LLM as Evaluator
2.3.1. LLM Evaluator Construction
We utilize pre-trained LLMs to provide annotations for each explanation text $E$. The large language model receives the movie name and the corresponding explanation text, accompanied by a prompt $\mathcal{P}$ to describe the context. Although $E$ is in Chinese as introduced in Section 2.2.1, we apply $\mathcal{P}$ in English to ensure a standardized prompt design. The LLM evaluation for each explanatory text $\mathbf{\hat{s}}_{LLM(u,i,E)}$ is a one-dimensional vector of length 4, with elements being integers between 1 to 5. Our experiments are based on item-wise evaluation, where only one text entry is evaluated at a time. This approach aligns with the setup in the user study.
@@ 2.3.1. LLM Evaluator Construction
Figure 2. The outline of evaluation prompt templates applied in our study.
2.3.2. Prompt Construction
@@
Single-Aspect vs. Multiple-Aspect. The difference between single and multiple-aspect prompts lies in the Aspect part. A multiple-aspect prompt instructs the model to evaluate all four aspects concurrently, while a single-aspect prompt focuses on one aspect at a time. Utilizing single-aspect prompts necessitates more times of interactions with LLM compared to multiple-aspect prompts. Experimental results under the two settings are presented and compared in RQ1, Section 2.4.
Algorithm 1 Comparison of personalized and non-personalized one-shot learning process
1:procedure PersonalizedOneShot
2:     for each user $u$ in $\mathcal{U}$ do
3:         Randomly choose item $i_{0}$ from $\mathcal{I}_{u}$
4:         Collect user evaluations on explanatory text $E=g(u,i_{0})$ as $\mathbf{s_{0}}_{g}$ for each $g$ in $\mathcal{G}$
5:         for each movie $i$ in $\mathcal{I}_{u}$, each $g$ in $\mathcal{G}$ do
6:              $Prompt\leftarrow$ ConstructPrompt($i$, $g(u,i)$, $\mathbf{s_{0}}_{g}$)
7:              $Result\leftarrow$ LLM($Prompt$)
8:              Parse $Result$ and process the answer
9:         end for
10:     end for
11:end procedure
12:procedure Non-PersonalizedOneShot
13:     Randomly choose user-item pair $(u_{0},i_{0})$
14:     Collect user evaluations on explanatory text $E=g(u_{0},i_{0})$ as $\mathbf{s_{0}}_{g}$ for each $g$ in $\mathcal{G}$
15:     for each user $u$ in $\mathcal{U}$,each movie $i$ in $\mathcal{I}_{u}$, each $g$ in $\mathcal{G}$ do
16:         $Prompt\leftarrow$ ConstructPrompt($i$, $g(u,i)$, $\mathbf{s_{0}}_{g}$)
17:         $Result\leftarrow$ LLM($Prompt$)
18:         Parse $Result$ and process the answer
19:     end for
20:end procedure
@@
To explore whether humans and LLM can collaborate on evaluation, we investigate the impact of utilizing human labels as contextual information in LLM evaluation. Our primary focus is on one-shot learning and includes one human label in the prompt. To enhance the LLM’s ability to learn personalized preferences, we employ personalized one-shot learning. This approach involves providing the scoring example from the same user for the target data. Formally, for the user $u$, item $i$ and explanation $E=g(u,i)$, the personalized example is the scoring given to $E^{\prime}=g(u,i_{0})$, which is the explanation generated by the same system $g$ and the same user $u$ on a randomly chosen movie item $i_{0}$. The methods help the prompt $\mathcal{P}$ incorporate personal information. However, a limitation is that it still requires collecting real user feedback, which is expansive and sometimes impractical. Hence, we also investigate non-personalized one-shot prompts. In this setting, all examples are from randomly selected $(u_{0},i_{0})$ pair. That is, the example can come from another user and are easier to collect in practice. Formally, for the user $u$, item $i$ and explanation $E=g(u,i)$, the non-personalized example is the scoring given to $E^{\prime}=g(u_{0},i_{0})$. We summarize the personalized and non-personalized one-shot learning procedures in Algorithm 1 to better illustrate the example selection process and their differences. Experimental results under different example settings are presented and compared in RQ2, Section 4.
2.4. Three-Level Meta Evaluation
Good evaluation requires that evaluation efforts themselves be evaluated (Stufflebeam et al., 1974). Employing the suitable meta-evaluation method $h(\cdot)$ is crucial for thoroughly examining the evaluation procedure. Correlation coefficients, such as Pearson($r$), Spearman($\rho$), and Kendall($\tau$) coefficients, have served as a widely used metric for gauging the similarity in trends between two arrays of ratings or scores (Liu et al., 2023b). In NLG, the strategy of meta-evaluation can be divided into two levels: Dataset-Level and Sample-Level (Liu et al., 2023b). Formally, given a dataset $\mathcal{D}$ consisting of a set of source data $\mathcal{X}=(x_{1},x_{2},....,x_{|\mathcal{X}|})$ and generation methods $\mathcal{G}=(g_{1}(\cdot),g_{2}(\cdot),...,g_{|\mathcal{G}|}(\cdot))$ and correlation metric $r(\cdot)$ dataset-level and sample-level meta evaluation are expressed as:
@@ 2.4. Three-Level Meta Evaluation
We use a concrete example to illustrate the possible variation of the same evaluation metric at different levels. For instance, our experimental results in Section 2.4 found that BLEU-4 scoring results correlate poorly or even negatively with ground-truth labels at the Dataset-Level, whereas correlations improve at User-Level and Pair-Level. This discrepancy arises because BLEU-4 computes the token similarity between the target text and the reference text. Some reference texts may contain more commonly used token combinations than others, potentially inflating the BLEU-4 score. This introduces a bias into the evaluation process that is unrelated to user perceptions, thereby impacting the Dataset-Level correlation. Full experimental results and analyses can be found in Section 2.4.
Table 1. The 3-level Pearson correlation between the results given by the evaluator and the ground-truth label given by users. Bold fonts denote the best results among all tested evaluators and the underlines show the second-best results.